# From Counterfactuals to Trees:Competitive Analysis of Model Extraction Attacks

### v1

@@ Abstract
The advent of Machine Learning as a Service (MLaaS) has heightened the trade-off between model explainability and security. In particular, explainability techniques, such as counterfactual explanations, inadvertently increase the risk of model extraction attacks, enabling unauthorized replication of proprietary models. In this paper, we formalize and characterize the risks and inherent complexity of model reconstruction, focusing on the “oracle” queries required for faithfully inferring the underlying prediction function. We present the first formal analysis of model extraction attacks through the lens of competitive analysis, establishing a foundational framework to evaluate their efficiency. Focusing on models based on additive decision trees (e.g., decision trees, gradient boosting, and random forests), we introduce novel reconstruction algorithms that achieve provably perfect fidelity while demonstrating strong anytime performance. Our framework provides theoretical bounds on the query complexity for extracting tree-based model, offering new insights into the security vulnerabilities of their deployment.
Machine learning, Model extraction attacks, Competitive analysis, Counterfactual explanations
@@ 1 Introduction
Recent research has shown that sharing trained machine learning (ML) models can lead to the reconstruction of sensitive training data, posing significant privacy risks (see, e.g., Boenisch et al., 2023; Carlini et al., 2024; Ferry et al., 2024). Applications in fields such as medical diagnostics, financial services, and personalized advertising often handle large amounts of private data, making them attractive targets for data reconstruction attacks. These attacks exploit vulnerabilities in the model to recover confidential information from the training dataset, thereby undermining the privacy guarantees that organizations seek to uphold. Consequently, organizations may prefer to utilize Machine Learning as a Service (MLaaS) to leverage powerful models without directly exposing them, balancing the benefits of advanced analytics with the need to protect sensitive information.
While MLaaS platforms provide accessible and scalable ML solutions, they must address the growing demand for explainability in their decision-making processes. Regulatory frameworks such as the EU AI Act’s Article 13111https://artificialintelligenceact.eu/article/13/ further mandate greater transparency across a wide range of applications. In response, MLaaS providers increasingly incorporate explainability techniques to elucidate model behavior and ensure fairness. However, studies have shown that querying a model’s explanations can enable attackers to replicate its hyperparameters and architecture, effectively copying the original model (Tramèr et al., 2016; Wang et al., 2022; Aïvodji et al., 2020; Öksüz et al., 2024). This reveals a critical tension between the need for transparency and the protection of model integrity and intellectual property.
Model extraction attacks were proposed against a variety of ML models in recent years (Oliynyk et al., 2023). While very few of them are functionally equivalent (i.e., they provably reconstruct the black-box model’s decision boundary), they often rely on strong assumption, such as access to a leaf identifier in the case of decision tree models (Tramèr et al., 2016). Moreover, the majority of the literature focuses solely on empirically evaluating the fidelity of the extracted model w.r.t. the target black-box, lacking a rigorous framework for analyzing attack complexities and thoroughly characterizing their worst-case scenarios. Finally, while counterfactual explanations constitute a promising attack surface and were exploited to conduct model extraction attacks (Aïvodji et al., 2020; Wang et al., 2022; Dissanayake & Dutta, 2024), existing approaches rely on training surrogate models without functional equivalence guarantees.
@@
Figure 1: Illustration of the connection between online discovery problems and model extraction attacks. The left figure was inspired by Figure 4 of Tee & Han (2021). In this context, an autonomous robot explores an unknown 2D environment (e.g., a house). The objective of the robot is to produce a map of this house, given a limited view of its surroundings provided by different sensors (e.g., LIDAR and laser distance measurements). The right figure corresponds to the problem tackled in this paper: model extraction attacks aim at recovering the decision boundary of the underlying model, by observing answers to counterfactual queries.
@@
  • We define a rigorous framework to characterize the complexity of model extraction attacks, utilizing competitive analysis (a notion from online optimization) to evaluate the difficulty of reconstructing models under various conditions
  • We introduce a novel algorithm (TRA) specifically designed to efficiently extract axis-parallel decision boundary models (including, but not limited to, tree ensemble models) through optimal counterfactual explanations.
  • We provide a comprehensive theoretical analysis of our proposed method, offering guarantees on query complexity and demonstrating 100% fidelity in the extracted models
  • We conduct extensive experiments to validate our theoretical findings, presenting an average-case and anytime performance analysis of TRA compared to state of the art reconstruction methods. These experiments not only confirm out theoretical results, but also provide practical insights into the effectiveness and limitations of our approach
@@ 2 Online Discovery, Model Extraction Attacks and Competitive Analysis
Online discovery problems have long been a focus of research in theoretical computer science, where the goal is to uncover the structure of an unknown environment through a sequence of queries or observations (Ghosh & Klein, 2010; Deng et al., 1991). A classic example arises in map exploration: an agent (e.g., a robot) navigates a space cluttered with obstacles, with only limited “line of sight” at each position. The agent’s objective is to construct a complete representation (e.g., map) of its surroundings while minimizing resources such as travel distance or exploration time.
@@
Online discovery problems & Model extraction attacks.
Online discovery typically assumes an agent that can move freely in the physical world while receiving feedback about obstacles in its vicinity. In model extraction, the “environment” is the model’s input space, and the queries return a point that lies on the nearest decision boundary (or provides the counterfactual boundary itself). Thus, while map exploration may allow richer geometric observations (e.g., an entire sensor sweep of obstacles), counterfactual-based model extraction often yields more constrained information (e.g., only the nearest boundary for a given input). Despite these differences, both problems share a common hallmark: the true structure (environment or decision boundaries) is unknown a priori and must be inferred online via carefully chosen queries.
Competitive Analysis.
A central tool for analyzing online problems is competitive analysis (Karlin et al., 1986), which compares the performance of an online algorithm — one that adapts its decisions based solely on information acquired so far — to an optimal offline algorithm that has full prior knowledge of the underlying structure. Formally, we measure the ratio between: (i) the complexity (e.g., number of queries, computational cost) incurred by the online algorithm to achieve its goal and (ii) the minimal complexity that an offline algorithm, with complete foresight, would require to accomplish the same task or verify its reconstruction. When this ratio remains bounded by a constant, we say that the online algorithm is constant-competitive. Otherwise, the ratio might grow with the size or complexity of the problem, revealing fundamental limitations of online approaches.
By applying competitive analysis to model extraction attacks, we link these two domains in a natural way. We quantify how many queries (i.e., counterfactual explanations or label predictions) are required to guarantee perfect fidelity in reconstructing a target model under worst-case scenarios. Moreover, competitive analysis encourages us to ask: How many queries, relative to an all-knowing attacker, does one need in order to prove with certainty that a specific model has been recovered? In this paper, we adapt these notions of worst-case online complexity to create a principled measure of the difficulty of extracting tree-based models, paralleling traditional results in online map discovery (Deng et al., 1991; Hoffmann et al., 2001; Ghosh & Klein, 2010; Fekete & Schmidt, 2010).
This perspective paves the way for a unified view: model extraction attacks can be seen as online exploration in feature space. Our work not only provides new theoretical results for tree-based model extraction but also invites cross-pollination between the literature on online discovery algorithms and emerging threats in machine learning security.
@@ 3.1 Problem Statement
We consider an input as an $m$-dimensionall vector in the input space $\mathcal{X}=\mathcal{X}_{1}\times\mathcal{X}_{2}\times\cdots\times\mathcal{X}_{m}\subseteq\mathbb{R}^{m}$, and the output belongs to the categorical space $\mathcal{Y}$. Let $\mathcal{F}$ denote the set of all axis-parallel decision boundary models, including decision trees and their ensembles, such as random forests (Breiman, 2001). A machine learning (ML) classification model $f\in\mathcal{F}$ is defined as a function $f:\mathcal{X}\mapsto\mathcal{Y}$. For simplicity and without loss of generality, we focus our discussion on decision trees, as any axis-parallel decision boundary model can be represented as a decision tree (Vidal & Schiffer, 2020). The input space $\mathcal{X}$ may comprise both categorical and continuous features.
@@
For a distance function $d$, an optimal counterfactual explanation oracle $\mathcal{O}_{d}$ is defined as a function $\mathcal{O}_{d}:\mathcal{F}\times\mathcal{X}\times\mathcal{P}(\mathcal{X})\mapsto\mathcal{X}$ that returns an optimal counterfactual explanation using the distance $d$ for a given instance $x\in\mathcal{X}$, a model $f\in\mathcal{F}$, and an input subset $\mathcal{E}\subseteq\mathcal{X}$:
$\mathcal{O}_{d}(f,x,\mathcal{E})\in\text{argmin}_{x^{\prime}\in\mathcal{E}}$ $\quad d(x,x^{\prime})$
s.t. $f(x^{\prime})\neq f(x)$
where $\mathcal{P}(\mathcal{X})$ is the set of all possible subsets $\mathcal{E}\subseteq\mathcal{X}$.
Formally, an optimal counterfactual explanation, also coined as closest counterfactual, can be interpreted as the nearest input to $x$ that yields a different prediction under model $f$.
For an adversary with black-box access to a target model $f$ (i.e., through a prediction API), the objective of a model extraction attack is to retrieve the exact model’s parameters (Tramèr et al., 2016). A tractable weakening, coined functionally equivalent extraction (Definition 3.2), is to reconstruct a model encoding the exact same function over the input space (Jagielski et al., 2020). Note that with such relaxation, the model’s parameters (or even hypothesis class) may differ as long as the prediction function is exactly the same. A common way to empirically evaluate such attacks is through the fidelity (Aïvodji et al., 2020) of the model reconstructed by the attacker, coined the surrogate model. It is defined as the proportion of examples (from a given dataset) for which the surrogate agrees with the target model.
@@
A functionally equivalent attack aims to reconstruct a model $\hat{f}\in\mathcal{F}$ such that it is functionally identical to the target model $f\in\mathcal{F}$ across the entire input space $\mathcal{X}$. Formally, the attack seeks to find $\hat{f}$ satisfying:
$\forall x\in\mathcal{X},\quad\hat{f}(x)=f(x)$
The objective is to achieve this equivalence with the minimum number of queries to the target model.
@@
Let $\mathcal{A}$ denote an online model extraction attack algorithm. Define $Q_{\mathcal{A}}^{f}$ as the number of queries required by $\mathcal{A}$ to extract the decision boundary model $f$, and let $Q_{opt}^{f}$ represent the minimal (optimal) number of queries necessary to extract $f$ by an omniscient offline algorithm. The algorithm $\mathcal{A}$ is said to be c-competitive if, for any model $f\in\mathcal{F}$:
@@
In this work, we address the problem of executing a functionally equivalent attack to reconstruct the decision function $f$ (i.e., decision boundary) using a minimal number of queries. To this end, our approach leverages access to an optimal counterfactual explanation oracle $\mathcal{O}_{d}$.
@@ 3.2 Tree Reconstruction Attack algorithm
We now introduce our primary algorithm, the Tree Reconstruction Attack (TRA), detailed in Algorithm 1. TRA is a divide-and-conquer based algorithm that aims to reconstruct a decision tree $f_{n}$ with $n$ split levels by leveraging a counterfactual explanation oracle $\mathcal{O}_{d}$ and systematically exploring the input space $\mathcal{X}$. A split level divides the input space into two subspaces based on a particular value for a given feature. Note that multiple nodes within different branches of a decision tree can share the same split level.
Algorithm Overview.
@@
TRA operates by maintaining a query list $\mathcal{Q}$ that initially contains the entire input space $\mathcal{X}$. The algorithm iteratively processes each input subset $\mathcal{E}\subseteq\mathcal{X}$ from $\mathcal{Q}$, until $\mathcal{Q}$ is empty, ensuring that all decision boundaries of the target model $f$ are identified and replicated in the reconstructed tree. More precisely, at each iteration, TRA first retrieves the subset $\mathcal{E}$ on top of the priority queue $\mathcal{Q}$. It computes its geometric center $x$ using the $center$ function (line 5). It then queries the oracle $\mathcal{O}_{d}$ with the current model $f$, input $x$, and subset $\mathcal{E}$ to obtain a counterfactual explanation $x^{\prime}=\mathcal{O}_{d}(f,x,\mathcal{E})$ (line 8). The set of feature indices where $x^{\prime}$ differs from $x$, i.e., $\{i\mid x^{\prime}_{i}\neq x_{i}\}$ is consequently identified. For each differing feature $i$, TRA splits the input subset $\mathcal{E}$ into two subspaces based on the split value $x^{\prime}_{i}$ ($split$ function). The resulting subspaces are added to $\mathcal{Q}$ for further exploration (line 9). If no counterfactual explanation $x^{\prime}$ exists within $\mathcal{E}$ , TRA assigns the label $y=f(x)$ to the entire subset $\mathcal{E}$, indicating that it corresponds to a leaf node in the reconstructed tree (line 11).
Illustrative Example.
@@
To demonstrate the execution of TRA, consider the axis-parallel decision boundary illustrated on the right side of Figure 2(b). Initially, TRA begins with the entire input space $\mathcal{X}=[0,1]^{2}$. In the first iteration, the algorithm selects the center point $x^{(1)}=(0.5,0.5)$ of $\mathcal{X}$ and queries the counterfactual explanation oracle $\mathcal{O}_{\lVert.\rVert_{2}}$, which returns a counterfactual $x^{\prime(1)}=(0.5,0.4)$ that differs from $x^{(1)}$ in the second feature ($x_{2}$). This results in the first split of the input space (to $\mathcal{E}_{1}=[0,1]\times]0.4,1]$ and $\mathcal{E}_{2}=\mathcal{X}\setminus\mathcal{E}_{1}$) based on the condition $x_{2}\leq 0.4$, as shown on the left side of Figure 2(b). In the subsequent iterations, TRA focuses on the resulting subspaces. For example, within the subset where $x_{2}\geq 0.4$, TRA identifies another split at $x_{1}\leq 0.7$, further partitioning the space. After three iterations, as shown in Figure 2(a), the reconstructed decision tree accurately captures part of the decision boundaries of the target model, effectively distinguishing between different regions in the input space. The gray hatched zones (or “?” nodes) represent regions that have not yet been explored and remain in the query list $\mathcal{Q}$.
@@
9:        $\mathcal{Q}\leftarrow\mathcal{Q}\cup split(\mathcal{E},x,x^{\prime})$Split $\mathcal{E}$ based on $x^{\prime}$ and add the resulting new subspaces to $\mathcal{Q}$}
@@
Let $f_{n}$ be a decision tree with $n$ split levels across a $m$-dimensional input space $\mathcal{X}=\mathcal{X}_{1}\times\mathcal{X}_{2}\times\cdots\times\mathcal{X}_{m}$. Denote $s_{i}$ as the number of split levels in $f_{n}$ over the $i$-th feature, such that $\sum_{i=1}^{m}s_{i}=n$. The worst-case complexity of Algorithm 1 is $O\left(\prod_{\begin{subarray}{c}i=1\\ s_{i}\neq 0\end{subarray}}^{m}s_{i}\right)$.
@@
Let $p\leq m$ denote the number of features with at least one split level. The worst-case complexity of Algorithm 1 is $O\left(\left({n}/{p}\right)^{p}\right)$.
The proofs of Proposition 3.4 and Corollary 3.5 are provided in the Appendix A. Proposition 3.4 establishes a first simple upper bound on the complexity of Algorithm 1, indicating that the algorithm’s efficiency is constrained by this theoretical limit. To illustrate, consider a two-dimensional decision tree that partitions the space in a chessboard-like pattern. A comprehensive mapping of such a space necessitates at least $s_{1}\times s_{2}$ queries (one in each sub-square), a requirement that holds for multi-class classification scenarios and in high-dimension ($m\geq 2$).
Query Selection Analysis.
The primary hyperparameter in the TRA algorithm, as outlined in Algorithm 1, pertains to the strategy for selecting query points within an input space. By default, TRA selects the geometrical center of the current input subset as the query point. Alternatively, one could choose other points such as the lower/upper left/right corners, or even a random point. In the following, we present theoretical results analyzing the impact of different query selection strategies on the algorithm’s performance. To this end, we leverage the notion of competitive analysis for online discovery problems discussed in Section 2.
@@
For $(n,m)\in\mathbb{N}^{2}$, the TRA Algorithm 1 achieves a competitive ratio of $C_{TRA}^{(n,m)}$, defined as:
$\leq\frac{2\left(1+\frac{n}{m}\right)^{m}-1}{2n+1},$
where $s_{i}$ is the number of splits along the $i$-th feature.

### v2

@@ Abstract
The advent of Machine Learning as a Service (MLaaS) has heightened the trade-off between model explainability and security. In particular, explainability techniques, such as counterfactual explanations, inadvertently increase the risk of model extraction attacks, enabling unauthorized replication of proprietary models. In this paper, we formalize and characterize the risks and inherent complexity of model reconstruction, focusing on the “oracle” queries required for faithfully inferring the underlying prediction function. We present the first formal analysis of model extraction attacks through the lens of competitive analysis, establishing a foundational framework to evaluate their efficiency. Focusing on models based on additive decision trees (e.g., decision trees, gradient boosting, and random forests), we introduce novel reconstruction algorithms that achieve provably perfect fidelity while demonstrating strong anytime performance. Our framework provides theoretical bounds on the query complexity for extracting tree-based models, offering new insights into the security vulnerabilities of their deployment.
@@ 1 Introduction
Recent research has shown that sharing trained machine learning (ML) models can lead to the reconstruction of sensitive training data, posing significant privacy risks [see, e.g., Boenisch et al., 2023, Carlini et al., 2024, Ferry et al., 2024]. Applications in fields such as medical diagnostics, financial services, and personalized advertising often handle large amounts of private data, making them attractive targets for data reconstruction attacks. These attacks exploit vulnerabilities in the model to recover confidential information from the training dataset, thereby undermining the privacy guarantees that organizations seek to uphold. Consequently, organizations may prefer to utilize Machine Learning as a Service (MLaaS) to leverage powerful models without directly exposing them, balancing the benefits of advanced analytics with the need to protect sensitive information.
While MLaaS platforms provide accessible and scalable ML solutions, they must address the growing demand for explainability in their decision-making processes. Regulatory frameworks such as the EU AI Act’s Article 13111https://artificialintelligenceact.eu/article/13/ further mandate greater transparency across a wide range of applications. In response, MLaaS providers increasingly incorporate explainability techniques to elucidate model behavior and ensure fairness. However, studies have shown that querying a model’s explanations can enable attackers to replicate its parameters and architecture, effectively copying the original model [Tramèr et al., 2016, Wang et al., 2022, Aïvodji et al., 2020, Öksüz et al., 2024]. This reveals a critical tension between the need for transparency and the protection of model integrity and intellectual property.
Model extraction attacks were proposed against a variety of ML models in recent years [Oliynyk et al., 2023]. While very few of them are functionally equivalent (i.e., they provably reconstruct the black-box model’s decision boundary), they often rely on strong assumptions, such as access to a leaf identifier in the case of decision tree models [Tramèr et al., 2016]. Moreover, the majority of the literature focuses solely on empirically evaluating the fidelity of the extracted model w.r.t. the target black-box, lacking a rigorous framework for analyzing attack complexities and thoroughly characterizing their worst-case scenarios. Finally, while counterfactual explanations constitute a promising attack surface and were exploited to conduct model extraction attacks [Aïvodji et al., 2020, Wang et al., 2022, Dissanayake and Dutta, 2024], existing approaches rely on training surrogate models without functional equivalence guarantees.
@@
Figure 1: Illustration of the connection between online discovery problems and model extraction attacks. Left (adapted from Tee and Han [2021]): an autonomous robot maps an unknown 2D environment (e.g., a house) with limited-range sensors (e.g., LIDAR and laser distance measurements). Right: model extraction attacks recover the model’s decision boundary via counterfactual queries.
@@
  • We define a rigorous framework to characterize the complexity of model extraction attacks, utilizing competitive analysis (a notion from online optimization) to evaluate the difficulty of reconstructing models under various conditions.
  • We introduce a novel algorithm (TRA) specifically designed to efficiently extract axis-parallel decision boundary models (including, but not limited to, tree ensemble models) through locally optimal counterfactual explanations.
  • We provide a comprehensive theoretical analysis of our proposed method, offering guarantees on query complexity and demonstrating 100% fidelity in the extracted models.
  • We conduct extensive experiments to validate our theoretical findings, presenting an average-case and anytime performance analysis of TRA compared to state of the art reconstruction methods. These experiments not only confirm our theoretical results, but also provide practical insights into the effectiveness and limitations of our approach.
@@ 2 Online Discovery, Model Extraction Attacks and Competitive Analysis
Online discovery problems have long been a focus of research in theoretical computer science, where the goal is to uncover the structure of an unknown environment through a sequence of queries or observations [Ghosh and Klein, 2010, Deng et al., 1991]. A classic example arises in map exploration: an agent (e.g., a robot) navigates a space cluttered with obstacles, with only a limited “line of sight” at each position. The agent’s objective is to construct a complete representation (e.g., map) of its surroundings while minimizing resources such as travel distance or exploration time.
@@
Online discovery problems & Model extraction attacks. Online discovery typically assumes an agent that can move freely in the physical world while receiving feedback about obstacles in its vicinity. In model extraction, the “environment” is the model’s input space, and the queries return a point that lies on the nearest decision boundary (or provides the counterfactual boundary itself). Thus, while map exploration may allow richer geometric observations (e.g., an entire sensor sweep of obstacles), counterfactual-based model extraction often yields more constrained information (e.g., only the nearest boundary for a given input). Despite these differences, both problems share a common hallmark: the true structure (environment or decision boundaries) is unknown a priori and must be inferred online via carefully chosen queries.
Competitive Analysis. A central tool for analyzing online discovery problems is competitive analysis [Karlin et al., 1986], which compares the performance of an online algorithm — one that adapts its decisions based solely on information acquired so far — to an optimal offline algorithm with complete foresight. Formally, we measure the ratio between: (i) the complexity (e.g., number of queries, computational cost) incurred by the online algorithm to achieve its goal and (ii) the minimal complexity that an offline algorithm, with complete foresight, would require to accomplish the same task or verify its reconstruction. A constant ratio implies a constant-competitive algorithm; in more complex settings, the ratio may grow with problem parameters.
By applying competitive analysis to model extraction attacks, we can quantify how many queries (i.e., counterfactuals or label predictions) are needed to guarantee perfect fidelity in model reconstruction under worst-case conditions, complementing empirical investigations. Moreover, competitive analysis encourages us to ask: How many queries, relative to an all-knowing attacker, does one need in order to prove with certainty that a specific model has been recovered? This yields a principled measure of the difficulty of extracting tree-based models, analogous to classic results in online map discovery [Deng et al., 1991, Hoffmann et al., 2001, Ghosh and Klein, 2010, Fekete and Schmidt, 2010].
Overall, this perspective paves the way for a unified view: model extraction attacks can be seen as online exploration in the feature space. Our work, therefore, provides new theoretical results for tree-based model extraction, and invites cross-pollination between the literature on online discovery algorithms and emerging threats in machine learning security.
@@ 3.1 Problem Statement
We consider an input as an $m$-dimensional vector in the input space $\mathcal{X}=\mathcal{X}_{1}\times\mathcal{X}_{2}\times\cdots\times\mathcal{X}_{m}\subseteq\mathbb{R}^{m}$, and the output belongs to the categorical space $\mathcal{Y}$. Let $\mathcal{F}$ denote the set of all axis-parallel decision boundary models, including decision trees and their ensembles, such as random forests [Breiman, 2001]. A machine learning (ML) classification model $f\in\mathcal{F}$ is defined as a function $f:\mathcal{X}\mapsto\mathcal{Y}$. For simplicity and without loss of generality, we focus our discussion on decision trees, as any axis-parallel decision boundary model can be represented as a decision tree [Vidal and Schiffer, 2020]. The input space $\mathcal{X}$ may comprise both categorical and continuous features.
@@
Let $d$ be a distance function and $\mathcal{P}(\mathcal{X})$ denote the set of all subsets of $\mathcal{X}$. A counterfactual explanation oracle $\mathcal{O}_{d}$ is defined as a function $\mathcal{O}_{d}:\mathcal{F}\times\mathcal{X}\times\mathcal{P}(\mathcal{X})\mapsto\mathcal{X}$. For a given model $f\in\mathcal{F}$, an instance $x\in\mathcal{X}$, and an input subspace $\mathcal{E}\subseteq\mathcal{X}$, $x^{\prime}=\mathcal{O}_{d}(f,x,\mathcal{E})$ is a counterfactual explanation such that $x^{\prime}\in\mathcal{E}$ and $f(x^{\prime})\neq f(x)$. This counterfactual is locally optimal if:
$\exists\,\epsilon>0\text{ such that }\forall\,v\in\mathbb{R}^{m}\setminus\{0\},\,\|v\|\leq\epsilon,\quad f(x)=f(x^{\prime}+v)\quad\text{or}\quad d(x,x^{\prime}+v)\geq d(x,x^{\prime}).$
Intuitively, this condition ensures that any small perturbation ($v$) of $x^{\prime}$ either gives an invalid counterfactual (having the same label as $x$) or increases the distance to the original input $x$.
For an adversary with black-box access to a target model $f$ (i.e., through a prediction API), a model extraction attack aims to retrieve the exact model’s parameters [Tramèr et al., 2016]. However, this goal is often too strict as many models might encode the same prediction function and thus remain indistinguishable through counterfactual or prediction queries. Consequently, a more tractable objective, known as functionally equivalent extraction (Definition 3.2), focuses on reconstructing a model encoding the exact same function over the input space [Jagielski et al., 2020].
@@
A functionally equivalent extraction attack aims to reconstruct a model $\hat{f}\in\mathcal{F}$ such that it is functionally identical to the target model $f\in\mathcal{F}$ across the entire input space $\mathcal{X}$. Formally, the attack seeks to find $\smash{\hat{f}}$ satisfying Equation (1) using as few queries as possible.
$\forall x\in\mathcal{X},\quad\hat{f}(x)=f(x)$ (1)
A common way to empirically evaluate such attacks is through the fidelity [Aïvodji et al., 2020] of the model reconstructed by the attacker, coined the surrogate model. It is defined as the proportion of examples (from a given dataset) for which the surrogate agrees with the target model. To theoretically bound the efficiency of a model extraction attack, we additionally rely on the notion of c-competitiveness from the online discovery literature, formalized in Definition 3.3.
@@
Let $\mathcal{A}$ denote an online model extraction attack algorithm. Define $Q_{\mathcal{A}}^{f}$ as the number of queries required by $\mathcal{A}$ to extract the decision boundary of model $f$, and let $Q_{opt}^{f}$ represent the minimal (optimal) number of queries necessary to extract $f$ by an omniscient offline algorithm. The algorithm $\mathcal{A}$ is said to be c-competitive if, for any model $f\in\mathcal{F}$:
@@
In this work, we focus on functionally equivalent model extraction attacks of axis-parallel decision boundary models. We assume that for each query $x$ to the API, the attacker obtains (i) the label $f(x)$ of the query and (ii) a locally optimal counterfactual explanation $x^{\prime}$.
@@ 3.2 Tree Reconstruction Attack algorithm
We now introduce our proposed Tree Reconstruction Attack (TRA), detailed in Algorithm 1. TRA is a divide-and-conquer based algorithm that aims to reconstruct a decision tree $f_{n}$ with $n$ split levels by systematically exploring the input space $\mathcal{X}$, using only a black-box API returning predictions and locally optimal counterfactuals. A split level is defined as a particular value for a given feature that divides the input space into two subspaces. Note that multiple nodes within different branches of a decision tree can share the same split level.
@@
Algorithm Overview. TRA operates by maintaining a query list $\mathcal{Q}$ that initially contains the entire input space $\mathcal{X}$. The algorithm iteratively processes each input subset $\mathcal{E}\subseteq\mathcal{X}$ from $\mathcal{Q}$, until $\mathcal{Q}$ is empty, ensuring that all decision boundaries of the target model $f$ are identified and replicated in the reconstructed tree. More precisely, at each iteration, TRA first retrieves the subset $\mathcal{E}$ on top of the priority queue $\mathcal{Q}$. It computes its geometric center $x$ using the $center$ function (line 5). It then queries the oracle $\mathcal{O}_{d}$ with the target model $f$, input $x$, and subset $\mathcal{E}$ to obtain a counterfactual explanation $x^{\prime}=\mathcal{O}_{d}(f,x,\mathcal{E})$ (line 8). The set of feature indices where $x^{\prime}$ differs from $x$, i.e., $\{i\mid x^{\prime}_{i}\neq x_{i}\}$ is consequently identified. For each differing feature $i$, TRA splits the input subset $\mathcal{E}$ into two subspaces based on the split value $x^{\prime}_{i}$ ($split$ function). The resulting subspaces are added to $\mathcal{Q}$ for further exploration (line 9). If no counterfactual explanation $x^{\prime}$ exists within $\mathcal{E}$ , TRA assigns the label $y=f(x)$ to $\mathcal{E}$, indicating that it corresponds to a leaf node in the reconstructed tree (line 11).
@@
Illustrative Example. To illustrate TRA’s operation, consider the axis-parallel decision boundary model illustrated on the right side of Figure 2(b). Initially, TRA begins with the entire input space $\mathcal{X}=[0,1]^{2}$. In the first iteration, the algorithm selects the center point $\smash{x^{(1)}}=(0.5,0.5)$ of $\mathcal{X}$ and queries the counterfactual explanation oracle $\smash{\mathcal{O}_{\lVert.\rVert_{2}}}$, which returns a counterfactual $\smash{x^{\prime(1)}}=(0.5,0.4)$ that differs from $\smash{x^{(1)}}$ in the second feature ($x_{2}$). This results in the first split of the input space (to $\mathcal{E}_{1}=[0,1]\times]0.4,1]$ and $\mathcal{E}_{2}=\mathcal{X}\setminus\mathcal{E}_{1}$) based on the condition $x_{2}\leq 0.4$, as shown on the left side of Figure 2(b). In the subsequent iterations, TRA focuses on the resulting subspaces. For example, within the subset where $x_{2}>0.4$, TRA identifies another split at $x_{1}\leq 0.7$, further partitioning the space. After three iterations, the reconstructed decision tree (shown in Figure 2(a)) accurately captures part of the decision boundaries of the target model, effectively distinguishing between different regions in the input space. The gray hatched zones (or “?” nodes) represent regions that have not yet been explored and remain in the query list $\mathcal{Q}$.
@@
9:        $\mathcal{Q}\leftarrow insert(split(\mathcal{E},x,x^{\prime}))$Split $\mathcal{E}$ based on $x^{\prime}$ and add the resulting subspaces to $\mathcal{Q}$}
@@
Let $f_{n}$ be a decision tree with $n$ split levels across a $m$-dimensional input space $\mathcal{X}=\mathcal{X}_{1}\times\mathcal{X}_{2}\times\cdots\times\mathcal{X}_{m}$. Denote $s_{i}$ as the number of split levels in $f_{n}$ over the $i$-th feature, such that $\sum_{i=1}^{m}s_{i}=n$. The worst-case complexity of Algorithm 1 is $O\left(\prod_{\begin{subarray}{c}i=1\end{subarray}}^{m}(s_{i}+1)\right)$.
@@
The worst-case complexity of Algorithm 1 is $O\left(\left(1+\frac{n}{m}\right)^{m}\right)$.
The proofs of Proposition 3.4 and Corollary 3.5 are provided in the Appendix A. Proposition 3.4 establishes a first simple upper bound on the complexity of Algorithm 1. Intuitively, consider a two-dimensional decision boundary that partitions the space in a chessboard-like pattern of size $s_{1}\times s_{2}$. A comprehensive mapping of such space necessitates at least $s_{1}\times s_{2}$ queries (one in each sub-square), a requirement that holds for multi-class classification scenarios and in high-dimension.
Query Selection Analysis. An important hyperparameter of TRA is the strategy used to select query points within the input space. By default, TRA selects the geometrical center of the current input subset $\mathcal{E}$ as the query point. Alternatively, one could choose other points such as the lower/upper left/right corners, or even a random point. In the following, we present theoretical results analyzing the impact of different query selection strategies on the algorithm’s performance. To this end, we leverage the notion of competitive analysis for online discovery problems discussed in Section 2.
@@
For $(n,m)\in\mathbb{N}^{2}$, Algorithm 1 achieves a competitive ratio of $C_{TRA}^{(n,m)}$, defined as:
$C_{TRA}^{(n,m)}=\frac{\sum_{i=1}^{m}s_{i}\prod_{j=1}^{i-1}(s_{j}+1)+\prod_{j=1}^{m}(s_{j}+1)}{2n+1}\leq\frac{2\left(1+\frac{n}{m}\right)^{m}-1}{2n+1},$
where $s_{i}$ is the number of split levels along the $i$-th feature within the tree $f_{n}$.